# Uncertainty-aware Reward Model: Teaching Reward Models to Know What is Unknown

### v1

@@ Abstract
Reward models (RM) play a critical role in aligning generations of large language models (LLM) to human expectations. However, prevailing RMs fail to capture the stochasticity within human preferences and cannot effectively evaluate the reliability of reward predictions. To address these issues, we propose Uncertain-aware RM (URM) and Uncertain-aware RM Ensemble (URME) to incorporate and manage uncertainty in reward modeling. URM can model the distribution of disentangled attributes within human preferences, while URME quantifies uncertainty through discrepancies in the ensemble, thereby identifying potential lack of knowledge during reward evaluation. Experiment results indicate that the proposed URM achieves state-of-the-art performance compared to models with the same size, demonstrating the effectiveness of modeling uncertainty within human preferences. Furthermore, empirical results show that through uncertainty quantification, URM and URME can identify unreliable predictions to improve the quality of reward evaluations.
@@ 1 Introduction
However, current RMs fail to capture the stochastic nature of human preferences (Baylis, 1950) and lack the ability to evaluate the reliability of the predicted rewards. In prevalent RMs, a value head (usually a linear layer) is added to the pretrained base model and maps the hidden states to reward scalars (Bai et al., 2022a; Ouyang et al., 2022) or attribute scores (Adler et al., 2024; Wang et al., 2024a). This results in a deterministic reward modeling process, unable to accommodate the variabilities of human preferences. Moreover, there is no other information to validate the reliability of these reward predictions.
Uncertainty is of major importance in machine learning (Hüllermeier & Waegeman, 2021) and an appropriate representation for uncertainty is essential for developing trustworthy and reliable models (Yang et al., 2009; Varshney & Alemzadeh, 2017). Uncertainty originates from two different sources: aleatoric and epistemic. Aleatoric uncertainty refers to the inherent variability and randomness of data. As opposed to this, epistemic uncertainty is caused by lack of knowledge, i.e. ignorance of the model instead of any underlying randomness.
In the context of reward modeling for LLM, aleatoric uncertainty refers to the stochasticity of human preferences, while epistemic uncertainty comes from the RMs’ lack of knowledge to make reliable evaluations. Therefore, introducing uncertainty to reward modeling improves modeling capacity of RMs and enhance reliability of the reward predictions. Consequently, by identifying and filtering out out-of-distribution (OOD) data where RMs fail to generalize, rewards with better reliability pave the way for more efficient alignments of LLMs.
In this paper, we propose Uncertain-aware RM (URM) and Uncertain-aware RM Ensemble (URME) to handle aleatoric and epistemic uncertainty in reward modeling respectively. URM is equipped with an uncertainty-aware value head to model the distributions of multiple attributes within human preferences. We demonstrate that with the popular bradley-terry-model loss function (Bradley & Terry, 1952), RMs cannot quantify uncertainty of human preferences even with an uncertainty-aware value head. Therefore, URMs are trained via maximum likelihood estimation and attribute regression. URME quantifies epistemic uncertainty by the discrepancies among URMs in the ensemble, identifying potential lack of knowledge. During reward evaluation, filtering strategy can be applied to prompt-response pairs with high uncertainty, in case that LLMs learn unintended or potentially harmful behaviors that URMs may not be able to accurately evaluate.
@@
Empirical results on a popular RM benchmark RewardBench (Lambert et al., 2024) demonstrate that URM with 8B model size achieves state-of-the-art performance among models with the same size and outperforms a number of strong large models including Nemotron-4-340B (Adler et al., 2024). And through uncertainty quantification, URM and URME are able to identify their level of knowledge for the input data and make the reward evaluations more reliable through filtering strategy. Furthermore, results of best-of-$n$ sampling validates that URM and URME can effectively enhance the generation quality of LLMs.
@@
(1) We introduce URM and URME to model the uncertainty within human preferences and reward models themselves.
(2) URM and URME are able to improve LLMs’ generation effectively. Notably, URM achieves state-of-the-art performance on RewardBench compared with models of the same size (8B).
(3) Empirical results demonstrate that URM and URME can successfully quantify uncertainty to identify areas where the models lack sufficient knowledge to make accurate predictions, leading to more reliable reward evaluations.
@@ 2 Preliminaries
LLM alignment typically consists of three stages (Ouyang et al., 2022): supervised fine-tuning (SFT), reward modeling and proximal policy optimization (PPO) (Schulman et al., 2017). SFT utilizes expert demonstrations to fine-tune the pretrained base model in a supervised-learning fashion to enable LLMs to follow user instructions.
Reward Modeling Reward modeling aims to learns human preferences explicitly (Ouyang et al., 2022) or implicitly (Rafailov et al., 2024). For some prompt $x$ and a response pair $(y_{w},y_{l})$, $y_{w}$ is the chosen response preferred by humans and $y_{l}$ is rejected. Following the Bradley-Terry model (Bradley & Terry, 1952), under RM $r_{\phi}$ the probability of $y_{w}$ being preferred than $y_{l}$, i.e. $y_{w}\succ y_{l}$, is
$P(y_{w}\succ y_{l}|x)$ $=\log\frac{\text{exp}(r_{\phi}(x,y_{w}))}{\text{exp}(r_{\phi}(x,y_{w}))+\text{exp}(r_{\phi}(x,y_{l}))}$ (1)
$=sigmoid(r_{\phi}(x,y_{w})-r_{\phi}(x,y_{l}))$
Thus, to train a RM to prioritize chosen responses over rejected responses, the loss function is the maximum likelihood estimation of Eq. 1
$L_{1}=-\mathbb{E}_{x,y_{w},y_{l}\sim D}\left[\log sigmoid\left(r_{\phi}(x,y_{w})-r_{\phi}(x,y_{l})\right)\right],$ (2)
where $r_{\phi}$ is the reward model parameterized by $\phi$, consisting of the pretrained base model and a linear value head. The trained RM can be used to improve the LLM’s generation by Best-of N (BoN) (Stiennon et al., 2020) or RLHF (Ouyang et al., 2022).
PPO In this stage, LLMs are fine-tuned with feedbacks from the RM. To prevent the model deviate too far from the pretrained model and forget linguistic skills, there is also an Kullback-Leibler (KL) divergence penality in the reward from the RM. Thus, the total reward $\hat{r}$ is
@@
where $\eta$ is the coefficient for the KL penalty, $\pi$ is the model to be fine-tuned and $\pi_{\text{ref}}$ is the reference model which is usually the SFT model. Running PPO to maximize reward from Eq. 3 can not only align the LLM with human preferences, but also prevent it from severely deviating from the reference model.
@@ 3.1 Multi-attribute Reward Modeling
To generate helpful, harmless and truthful responses (Askell et al., 2021), LLMs must be aligned with human expectations. Current methods fine-tune models based on human (Christiano et al., 2017; Stiennon et al., 2020; Bai et al., 2022a; Ouyang et al., 2022) or AI feedbacks (Bai et al., 2022b; Sun et al., 2024) to maximize preference-based rewards, which are provided by reward modeling. Typically, a reward model is learned and LLMs will improve their generation quality according to feedbacks from the reward model (Bai et al., 2022a; Ouyang et al., 2022; Shao et al., 2024; Stiennon et al., 2020).
@@
Recent studies show that human and LLM judges may introduce potential biases to annotations of preference (Zhang et al., 2023; Kotek et al., 2023; Wang et al., 2024b; Chen et al., 2024a). Moreover, traditional RMs usually rely on single-dimensional feedback on general quality instead of fine-grained multifaceted signals to indicate multiple attributes such as helpfulness, coherence and verbosity (Dong et al., 2023b). Adler et al. (2024) discovered that multi-attribute RMs trained on datasets with high-quality attribute-specific annotations (Cui et al., 2023; Wang et al., 2024c) are able to disentangle real helpfulness and other irrelevant aspects such as lengthy bias (Shen et al., 2023; Singhal et al., 2023b). There are also alignment methods directly aimed at multi-attribute alignment. Zhou et al. (2023) includes preference on multiple attributes in the Direct Preference Optimization (DPO) loss function (Rafailov et al., 2024), trying to optimize preference rewards for all attributes simultaneously. Lou et al. (2024) proposed to achieve multi-attribute alignment sequentially, one attribute at a time, where LLMs learns to align with new attributes while staying aligned with previous dimensions.
3.2 RLHF, Offline RL and Uncertainty
@@
In RLHF, LLM policy is optimized via interactions with the RM, whose training data is pre-collected preference pairs (Bai et al., 2022a; Ouyang et al., 2022). In this setting, RLHF falls into the category of offline RL, where RL policies cannot interact with the environment and get feedbacks in real time, but instead can only be updated based on an offline dataset collected by some other policy (Levine et al., 2020). Offline RL is notoriously difficult due to the distributional shift issue (Lou et al., 2022; Ma et al., 2021; Prudencio et al., 2023). Recent advancements in iterative LLM alignment methods (Yuan et al., 2024; Dong et al., 2024; Xiong et al., 2024) iterates between LLM fine-tuning and the sampling and annotation of new training data, alleviating the distributional shift issue. Although these iterative methods aim to transcend the constraints of the offline setting,, RLHF is still offline within each iteration.
An important topic in offline RL is uncertainty quantification, which has long underpinned many critical roles (Abdar et al., 2021), such as trustworthy decision-making (Huang et al., 2019; Eriksson & Dimitrakakis, 2019) and improving reliability of machine learning models (Wang et al., 2019). In offline RL, uncertainty quantification enables out-of-distribution data detection and keep the policy within the offline dataset’s support area through conservative updates to avoid distributional shift (Yu et al., 2020; Kidambi et al., 2020; An et al., 2021; Zhu et al., 2024). So it is natural to introduce uncertainty to RLHF to make LLM alignment more reliable and effective.
@@
Ensemble of RMs are discussed in previous works. However, we study the ensemble of uncertainty-aware RMs to identify unreliable reward evaluations, while previous discussions are limited to using RM ensembles to mitigate reward hacking Coste et al. (2023); Eisenstein et al. (2023) and using value heads to disentangle length and quality in reward modeling Chen et al. (2024b). We notice a concurrent work QRM Dorka (2024) which also models human preferences by distributions. QRM only studies distributional RMs, while we also study the ensemble of such uncertainty-aware RMs. Moreover, QRM is trained via quantile regression (Koenker, 2017), a variant of our attribute regression. But we also studied uncertainty-aware RMs trained via maximum likelihood estimation, which can better capture the uncertainty of rewards.
@@ 4 Methodology
Figure 1: Architecture of URM and URME. URMs output $\mu$ and $\sigma$ to parameterize normal distributions, from which multi-attribute scores are sampled. The scores are then combined to reward scalars by weights generated by a gating layer. URME consists of multiple URMs, allowing for quantification of the epistemic uncertainty using the disagreement among the URMs. In the given example, there is a substantial divergence among URMs, indicating significant epistemic uncertainty. Thus, although URM 1 correctly assigns a small, negative reward to the input, the significant epistemic uncertainty still indicates the URMs lack relevant knowledge to provide reliable evaluation of the inputs.
@@ 4.1 Uncertainty-aware Reward Model
Traditional RMs optimize the Bradley-Terry model loss (BT-loss) in Eq. 2,to enlarge the discrepancy between the rewards of chosen and rejected responses so that the preference probability is maximized. The value head will map hidden states from the base model to a scalar reward. Such mapping is deterministic and thus cannot catch any uncertainty (Chua et al., 2018) within the reward modeling process.
@@
However, at its core, human preferences exhibit a distinctly probabilistic nature, rather than being strictly deterministic (Baylis, 1950). This issue is further exaggerated due to the bias and inconsistencies introduced by human annotators (Sylolypavan et al., 2023; Sleeman & Gilhooly, 2023; Chen et al., 2024a). Between individuals, preferences differ from person to person. This means what’s preferable for one may not be for another. Even within individuals, preferences are not static. They can swing based on numerous factors such as mood and context. These stochastic natures of human preferences contribute to adopting a probabilistic framework for modeling preferences with aleatoric uncertainty.
Prior works have explored a number of uncertainty-aware neural networks (Neal, 2012; Lakshminarayanan et al., 2017), especially in RL (Gal et al., 2016; Depeweg et al., 2016) and model-based RL (MBRL) (Chua et al., 2018; Yu et al., 2020; Kidambi et al., 2020).
Figure 2: Comparison between URM and traditional RM in estimating preference distribution.
@@
Considering RMs act similarly to the reward part of the dynamics model in MBRL, aleatoric uncertainty within human preferences can be captured by outputting the parameters of a parameterized distribution. Specifically, unlike traditional RMs that output a single deterministic reward value, uncertainty-aware RMs can model the distributions of human preferences. As schematically shown in Fig. 2, given a prompt-response pair with multiple preference annotation samples, traditional RMs can only provide a fixed reward estimation and fails to represent the real preference. But uncertainty-aware RMs are able to offer a more accurate approximation of the human preference distribution.
@@
To model the preference reward distribution, URM adds an probabilistic value head to the pretrained base model. The value head takes in the last hidden state $h$ of the base model and outputs mean $\mu$ and logged standard deviation $\sigma$ to parameterize a normal distribution $\mathcal{N}(\mu,\text{exp}(2\sigma))$, from which preference rewards are sampled, and the aleatoric uncertainty is quantified by variance of the distribution. Reparameterization technique is adopted to enable gradient back-propagation.
However, we show that introducing the probabilistic value head and the sample-based reward to RMs with BT-loss, the aleatoric uncertainty still cannot be quantified.
We denote the reparameterization parameter $\alpha\sim\mathcal{N}(0,1)$, and thus the sampled reward $r=\mu+\alpha\text{exp}(\sigma)$. Substituting chosen reward $r_{w}$ and rejected reward $r_{l}$ into BT-loss in Eq. 2. For some given input $x,y_{w},y_{l}$, gradient w.r.t. the logged standard deviation $\sigma_{w}$ is given by
$\nabla_{\sigma_{w}}L_{1}=-\mathbb{E}_{\alpha_{w},\alpha_{l}\sim\mathcal{N}(0,1)}\left[\alpha_{w}\big{(}1-sigmoid(r_{w}-r_{l})\big{)}\text{exp}(\sigma_{w})\right]=0,$
where $r_{w}=\mu_{w}+\alpha_{w}\text{exp}(\sigma_{w})$ and $r_{l}=\mu_{l}+\alpha_{l}\text{exp}(\sigma_{l})$. Similarly, $\nabla_{\sigma_{l}}L_{1}=0$. The unlearning effect of the variance term demonstrates that under the BT-loss, RMs still cannot quantify the uncertainty even equipped with a probabilistic value head.
Recent advances in multi-attribute RMs demonstrate they are capable of providing fine-grained rewards and disentangling real helpfulness and other irrelevant aspects such as lengthy bias (Adler et al., 2024; Chen et al., 2024a). The multi-attribute scores consist of human-or-AI-annotated ratings on multiple aspects such as helpfulness, coherence and verbosity. To learn a multi-attribute uncertainty-aware RM, we propose two ways to train the probabilistic value head.
@@
$L_{2}=-\mathbb{E}_{x,y\sim D}$ $\left[\log\mathbf{P_{\theta}}(R|x,y)\right]=-\mathbb{E}_{x,y\sim D}\left[\sum\limits_{i=0}^{n}\log P_{\theta}(R_{i}|x,y)\right],$ (4)
@@
[…]
With the trained probabilistic value head, during inference we can use mean $\mu_{i}$ for each attribute $i$ as the scores. We learn a gating layer to combine the multi-attribute scores to a reward scalar via weighted sum (Wang et al., 2024a). The gating layer is a fully-connected network, whose input is the last hidden states of the base LLM. And the learning objective of the gating layer is to prioritize chosen responses over rejected responses through the BT loss. For some prompt $x$, chosen response $y_{w}$ and rejected response $y_{l}$, the gating layer will output weights $\omega$ to combine the scores and thus the loss function is given by
$L_{4}=-\mathbb{E}_{x,y_{w},y_{l}\sim D}\left[\log sigmoid\left(\mu^{T}(x,y_{w})\omega(x,y_{w})-\mu^{T}(x,y_{l})\omega(x,y_{l})\right)\right],$ (6)
Since the gating layer only offers weights to combine the scores, the base model and the probabilistic value head are kept frozen during training the gating layer. Besides the gating layer, the multi-attribute scores can also be combined by prior weights and still demonstrate competitive performance (Adler et al., 2024).

### v2

@@ Abstract
Reward models (RMs) are essential for aligning large language models (LLM) with human expectations. However, existing RMs struggle to capture the stochastic and uncertain nature of human preferences and fail to assess the reliability of reward predictions. To address these challenges, we introduce the Uncertainty-aware Reward Model (URM) and its ensemble variant, URME. URM employs a probabilistic value head to capture aleatoric uncertainty by modeling the distribution of disentangled human preference attributes. URME further quantifies epistemic uncertainty by examining discrepancies among individual URMs within the ensemble, enabling identification of unreliable evaluations. Our empirical evaluations demonstrate that URM achieves strong performance on RewardBench, outperforming competitive large-scale models. Additionally, extensive experiments—including best-of-n sampling (BoN), iterative direct preference optimization (iterative DPO), and proximal policy optimization (PPO)—demonstrate that URM and URME significantly enhance LLMs’ generation quality. Notably, reward predictions with lower uncertainty are far more reliable, demonstrate significantly higher quality, and result in substantially improved alignment.
Uncertainty-aware Reward Model: Teaching Reward Models to Know What is Unknown
Xingzhou Lou1,2, Dong Yan3, Wei Shen3, Yuzi Yan4, Jian Xie3, Junge Zhang1,2,thanks: Correspondence. 1Institute of Automation, Chinese Academy of Sciences 2School of Artificial Intelligence, University of Chinese Academy of Sciences 3Baichuan Inc. 4Department of Electronic Engineering, Tsinghua University louxingzhou2020@ia.ac.cn, {sproblvem, xiejian1990}@gmail.com, shenwei@baichuan-inc.com,yan-yz17@tsinghua.org.cn, jgzhang@nlpr.ia.ac.cn
@@ 1 Introduction
Figure 1: Comparison between URM and traditional RM in estimating preference reward distribution.
Human preferences are inherently probabilistic rather than strictly deterministic (Baylis, 1950), yet current RMs fail to account for this stochasticity. As illustrated in Fig. 1, traditional RMs provide point-wise reward estimates that fall short of capturing the true reward distribution. Furthermore, these models are also unable to evaluate the reliability of their predictions. Consequently, inaccurate reward signals can misalign LLMs, resulting in suboptimal or even harmful behavior due to unreliable or biased feedback.
Proper representation of uncertainty is essential for developing reliable machine learning models (Hüllermeier and Waegeman, 2021; Yang et al., 2009; Varshney and Alemzadeh, 2017). Uncertainty arises from two sources: aleatoric, due to inherent data randomness, and epistemic, caused by model ignorance and lack of knowledge. In reward modeling for LLMs, aleatoric uncertainty reflects the stochasticity of human preferences, while epistemic uncertainty stems from the RM’s lack of knowledge for accurate evaluation. Incorporating uncertainty into reward modeling enhances RM’s modeling capacity and reliability by identifying and filtering out out-of-distribution (OOD) and untrustworthy reward predictions, thereby improving the alignment of LLMs.
In this paper, we propose the Uncertainty-aware Reward Model (URM) and its ensemble variant, URME, to address aleatoric and epistemic uncertainty in reward modeling. URM is equipped with an uncertainty-aware value head that outputs a parameterized distribution to model multiple attributes within human preferences. URME quantifies epistemic uncertainty by measuring discrepancies among the individual URMs in the ensemble, identifying areas of insufficient knowledge. Although uncertainty of the two sources is usually hard to distinguish (Hüllermeier and Waegeman, 2021), high uncertainty of either source indicates the model’s inability to make confident and reliable reward predictions. Thus, during reward evaluation and preference optimization, data with high uncertainty can be identified and filtered out to prevent LLMs from learning unintended or potentially harmful behaviors.
@@
Our empirical experiments consist of evaluations for both reward modeling performance and LLM alignment. For evaluation of reward modeling, results on RewardBench (Lambert et al., 2024) demonstrate that URM with 8B model size achieves strong performance among models with the same size and outperforms a number of strong large models including Nemotron-4-340B (Adler et al., 2024). And through uncertainty quantification, both URM and URME are able to identify unreliable reward predictions with high uncertainty and provide more trustworthy reward signals through filtering.
For evaluation of LLM alignment with URM, we conduct experiments using Best-of-N (BoN) sampling, iterative DPO Xiong et al. (2024) and RLHFOuyang et al. (2022) on AlpacaEval Li et al. (2023). Results of BoN validates that URM and URME can effectively enhance the generation quality of LLMs. And experiments on iterative DPO and RLHF indicate that compared to using all data indiscriminately, leveraging reward predictions with low uncertainty significantly improves reward quality and leads to substantially stronger performance in aligning LLMs.
@@
(1) We introduce URM and URME to model and quantify uncertainty within human preferences and reward models.
(2) The proposed models achieve strong reward modeling performance, surpassing baselines and competitive large-scale models
(3) Experiments across multiple methods show that uncertainty quantification enables URM to provide reliable reward predictions, significantly improving LLM alignment and generation quality.
@@ 2 Preliminaries
LLM alignment typically consists of three stages (Ouyang et al., 2022): supervised fine-tuning (SFT), reward modeling and preference optimization. SFT utilizes expert demonstrations to fine-tune the pretrained model to enable LLMs to follow user instructions.
Reward Modeling Reward modeling aims to learns human preferences explicitly (Ouyang et al., 2022) or implicitly (Rafailov et al., 2024). For prompt $x$ and a response pair $(y_{w},y_{l})$, $y_{w}$ is the chosen response preferred by humans and $y_{l}$ is rejected. The aim of reward modeling is to prioritize chosen responses over rejected responses, i.e. $y_{w}\succ y_{l}$, where the order is determined by BT model (Bradley and Terry, 1952) in pairwise ranking RMs (Ouyang et al., 2022) and specific attribute scores (Cui et al., 2023; Wang et al., 2024c) in multi-attribute RMs (Adler et al., 2024).
Preference Optimizaztion In this stage, LLMs are trained with feedbacks from the RM. RLHF fine-tunes the SFT model $\pi$ with the following reward
@@
where the KL penalty prevents the model from severe deviation and $\pi_{\text{ref}}$ is the reference model. DPO (Rafailov et al., 2024) directly fine-tunes the model to prefer the chosen response over the rejected via a simple classification loss. Iterative DPO (Xiong et al., 2024) extends DPO to the online setting, where DPO are iteratively applied with alternating stages of model fine-tuning and data collection. This iterative process is able to progressively enhance LLMs’ alignment.
@@ 3.1 Multi-attribute Reward Modeling
To generate helpful, harmless and truthful responses (Askell et al., 2021), LLMs must be aligned with human expectations. Current methods fine-tune models based on human (Christiano et al., 2017; Stiennon et al., 2020; Bai et al., 2022a; Ouyang et al., 2022) or AI feedbacks (Bai et al., 2022b; Sun et al., 2024) to maximize preference-based rewards, which are provided by RMs.
@@
Recent studies show that human and LLM judges may introduce biases to annotations of preference (Zhang et al., 2023; Kotek et al., 2023; Wang et al., 2024b; Chen et al., 2024a). Moreover, traditional RMs usually rely on single-dimensional feedback on general quality instead of fine-grained multifaceted signals to indicate multiple attributes such as helpfulness and coherence (Dong et al., 2023b). Adler et al. (2024) discovered that multi-attribute RMs trained on datasets with high-quality attribute-specific annotations (Cui et al., 2023; Wang et al., 2024c) are able to disentangle real helpfulness and other irrelevant aspects such as lengthy bias (Shen et al., 2023; Singhal et al., 2023b).
3.2 RLHF and Uncertainty
@@
In RLHF, LLM policy is optimized via interactions with the RM, whose training data is pre-collected (Bai et al., 2022a; Ouyang et al., 2022). In this setting, RLHF falls into the category of offline RL, where RL policies cannot interact with the environment and get feedbacks in real time, but instead can only be updated based on an offline dataset collected by some other policy (Levine et al., 2020). Offline RL is notoriously difficult due to the distributional shift issue (Lou et al., 2022; Ma et al., 2021; Prudencio et al., 2023). Recent advancements in iterative LLM alignment methods (Yuan et al., 2024; Dong et al., 2024; Xiong et al., 2024) iterates between LLM fine-tuning and the sampling and annotation of new training data, alleviating the distributional shift issue. Although these iterative methods aim to transcend the constraints of the offline setting,, RLHF is still offline within each iteration. In offline RL, uncertainty quantification enables OOD data detection and keep the policy within the offline dataset’s support area through conservative updates to avoid distributional shift (Yu et al., 2020; Kidambi et al., 2020; An et al., 2021; Zhu et al., 2024). So it is natural to introduce uncertainty to RLHF to make LLM alignment more reliable and effective.
@@
Ensemble of RMs are discussed in previous works. However, we study the ensemble of uncertainty-aware RMs to identify unreliable reward predictions, while previous discussions are limited to using RM ensembles to mitigate reward hacking Coste et al. (2023); Eisenstein et al. (2023) and using value heads to disentangle length and quality in reward modeling (Chen et al., 2024b). We notice a concurrent work QRM (Dorka, 2024) which also models preferences by distributions. QRM only studies distributional RMs, while we also study the ensemble of such uncertainty-aware RMs. Moreover, QRM is trained via quantile regression (Koenker, 2017), a variant of our attribute regression. But we also studied uncertainty-aware RMs trained via maximum likelihood estimation, which can better capture the uncertainty of rewards.
@@ 4 Methodology
Figure 2: Architecture of URM and URME. URMs output $\mu$ and $\sigma$ to parameterize normal distributions, from which multi-attribute scores are sampled. The scores are then combined to reward scalars by pre-determined weights, which can also be generated by a gating layer. URME consists of multiple URMs, allowing for quantification of the epistemic uncertainty using the disagreement among the URMs. In the given example, there is a substantial divergence among URMs, indicating significant epistemic uncertainty, which indicates the URMs lack relevant knowledge to provide reliable evaluation of the inputs.
@@ 4.1 Uncertainty-aware Reward Model
Recent advances demonstrate that compared to BT-model RMs Ouyang et al. (2022); Bai et al. (2022a), multi-attribute RMs can provide fine-grained, steerable rewards and effectively disentangle true helpfulness from irrelevant aspects such as length bias (Adler et al., 2024; Chen et al., 2024a). Multi-attribute scores are based on human or AI-annotated ratings across various aspects like helpfulness, coherence, and verbosity. The steerable rewards in multi-attribute RMs combined with uncertainty quantification holds the promise of building more reliable and controllable LLM evaluators. Consequently, we develop URMs that account for multiple attributes in human preferences.
@@
Value heads of traditional RMs will map hidden states from the base model to a scalar reward. Such mapping is deterministic and thus cannot catch any uncertainty (Chua et al., 2018) within the reward modeling process.
@@
However, at its core, human preferences exhibit a distinctly probabilistic nature, rather than being strictly deterministic (Baylis, 1950). This issue is further exaggerated due to the bias and inconsistencies introduced by human annotators (Sylolypavan et al., 2023; Sleeman and Gilhooly, 2023; Chen et al., 2024a). Between individuals, preferences differ from person to person. This means what’s preferable for one may not be for another. Even within individuals, preferences are not static. They can swing based on numerous factors such as mood and context. These stochastic natures of human preferences contribute to adopting a probabilistic framework for modeling preferences with aleatoric uncertainty.
@@
To capture the aleatoric uncertainty within human preferences, we can use a parameterized distribution to represent the preference rewards. Specifically, unlike traditional RMs that output a single deterministic reward value, uncertainty-aware RMs can model the distributions of human preferences. As schematically shown in Fig. 1, given a prompt-response pair with multiple preference annotation samples, traditional RMs can only provide a fixed reward estimation and fails to represent the real preference. But uncertainty-aware RMs are able to offer a more accurate approximation of the human preference distribution.
@@
To model the preference reward distribution, URM adds an probabilistic value head to the pretrained base model. The value head takes in the last hidden state $h$ of the base model and outputs mean $\mu$ and logged standard deviation $\sigma$ to parameterize a normal distribution $\mathcal{N}(\mu,\text{exp}(2\sigma))$, from which preference rewards are sampled, and the aleatoric uncertainty is quantified by variance of the distribution. To learn a multi-attribute uncertainty-aware RM, we propose two ways to train the probabilistic value head.
@@
$L_{2}$ $=-\mathbb{E}_{x,y\sim D}\left[\log\mathbf{P_{\theta}}(R|x,y)\right]$ (2)
$=-\mathbb{E}_{x,y\sim D}\left[\sum\limits_{i=0}^{n}\log P_{\theta}(R_{i}|x,y)\right],$
@@
where $i$ indicates $i$-th attribute, and $r_{i}\sim\mathcal{N}\left(\mu_{i},\text{exp}(2\sigma_{i})\right)$ is sampled from the distribution parameterized by the output of the probabilistic value head. To enable gradient back-propagation, we use the reparameterization technique, so that $r=\mu+\alpha\text{exp}(\sigma)$, where reparameterization parameter $\alpha\sim\mathcal{N}(0,1)$. A more detailed analysis of this MSE loss is given in the appendix B.
With the trained probabilistic value head, we can combine the multi-attribute scores to a reward scalar via weighted sum where weights are pre-determined (Adler et al., 2024) or generated by a gating layer (Wang et al., 2024a). In our multi-attribute URMs, the estimated aleatoric uncertainty is the sum of variances of the attributes.